### v3

# BlockPruner: Fine-grained Pruning for Large Language Models

Abstract
With the rapid growth in the size and complexity of large language models (LLMs), the costs associated with their training and inference have escalated significantly. Research indicates that certain layers in LLMs harbor substantial redundancy, and pruning these layers has minimal impact on the overall performance. While various layer pruning methods have been developed based on this insight, they generally overlook the finer-grained redundancies within the layers themselves. In this paper, we delve deeper into the architecture of LLMs and demonstrate that finer-grained pruning can be achieved by targeting redundancies in multi-head attention (MHA) and multi-layer perceptron (MLP) blocks. We propose a novel, training-free structured pruning approach called BlockPruner. Unlike existing layer pruning methods, BlockPruner segments each Transformer layer into MHA and MLP blocks. It then assesses the importance of these blocks using perplexity measures and applies a heuristic search for iterative pruning. We applied BlockPruner to LLMs of various sizes and architectures and validated its performance across a wide range of downstream tasks. Experimental results show that BlockPruner achieves more granular and effective pruning compared to state-of-the-art baselines. Our code is available at: https://github.com/MrGGLS/BlockPruner.
BlockPruner: Fine-grained Pruning for Large Language Models
Longguang Zhong1, Fanqi Wan1, Ruijun Chen1, Xiaojun Quanthanks: Corresponding author1, Liangzhi Li2 1School of Computer Science and Engineering, Sun Yat-sen University 2Meetyou AI Lab {zhonglg5,wanfq,chenrj8}@mail2.sysu.edu.cn quanxj3@mail.sysu.edu.cn, liliangzhi@xiaoyouzi.com
1 Introduction
Large language models (LLMs) (Zhao et al., 2023; Minaee et al., 2024) have demonstrated outstanding performance across a diverse array of natural language processing tasks. However, their growing size and complexity have led to substantial computational demands and increased memory usage, creating obstacles for deployment in resource-constrained environments. Model compression techniques (Gao et al., 2020; Li et al., 2023; Wang et al., 2024) have emerged as a promising solution to address the challenges of deploying large, computationally intensive models. These techniques aim to transform large models into more compact versions that require less storage and execute with lower latency, while minimizing performance degradation. Model compression methods typically involve knowledge distillation (Huang et al., 2022; Gu et al., 2024), quantization (Yao et al., 2022; Dettmers et al., 2023), and pruning (van der Ouderaa et al., 2024; Ashkboos et al., 2024). In this study, we primarily focus on pruning, a technique that can be combined with these other methods to achieve more effective and efficient compression.
Figure 1: Block Influence (BI) scores (Men et al., 2024) for the Llama2-7B model (Touvron et al., 2023b) computed at both layer and block levels, where blocks/layers with lower BI scores indicate less importance. The model has 32 Transformer layers, each containing one MHA and one MLP block, totaling 64 blocks. Block-level BI scores are generally lower than layer-level scores, indicating finer-grained redundancies.
Recent research on layer redundancy has shown that LLMs contain a substantial number of redundant layers (Yang et al., 2024; Men et al., 2024; Chen et al., 2024). Removing these layers does not severely impact the model’s performance. To quantify this redundancy, researchers have investigated various similarity-based measurement methods and developed corresponding pruning strategies, including layer merging (Yang et al., 2024) and layer removal (Men et al., 2024). These methods not only maintain the original width of the model architecture and avoid introducing additional structures, but also demonstrate superior performance. Furthermore, Gromov et al. (2024) posited that this observed redundancy may be intrinsically linked to the residual structure (He et al., 2016) inherent in the Transformer architecture. Building on this intuition and recognizing that Transformer layers can be further subdivided into smaller residual blocks, namely multi-head attention (MHA) and multi-layer perceptron (MLP) [^1] In this work, unless otherwise specified, we refer to a block as one of the two sublayers: MHA or MLP., we hypothesize that fine-grained block redundancies could exist within LLMs. Consequently, we conducted a preliminary experiment to assess the significance of blocks at varying granularities. Specifically, we sampled 32 instances from the Alpaca dataset (Taori et al., 2023) and employed the Block Influence (BI) metric (Men et al., 2024) to evaluate blocks at layer and block levels, as depicted in Figure 1. The results reveal that block-level BI scores are generally lower than layer-level BI scores, indicating that fine-grained redundancies at the block level are more significant within the model.
Building on these findings, we argue that finer-grained pruning can be effectively implemented in LLMs. Therefore, we introduce BlockPruner, a novel, training-free structured pruning approach. Unlike existing methods that focus on entire layers, BlockPruner segments each Transformer layer into multi-head attention (MHA) and multi-layer perceptron (MLP) blocks. It then evaluates the importance of these blocks using perplexity measures and applies a heuristic search for iterative pruning.
To validate the effectiveness of our method, we applied BlockPruner to six LLMs of varying sizes and architectures, and evaluated their performance using five representative benchmarks. Our experimental results demonstrate that BlockPruner provides more granular and effective pruning compared to state-of-the-art baselines. Additionally, we performed a series of analytical experiments to investigate the impact of block type, block importance metrics, and data on pruning effectiveness. Our findings confirm that LLMs contain substantial redundancies at the block level compared to the layer level, demonstrating that fine-grained pruning is more effective and appropriate than layer-based approaches for compressing these models.
2 Related Work
Pruning is a well-established technique to compress and accelerate neural networks by removing superfluous weights or structures within models. Pruning methods can be broadly categorized into unstructured pruning and structured pruning.
Unstructured pruning.
Unstructured pruning targets individual weights, eliminating redundant connections in neural networks by setting the corresponding weights to zero. For instance, SparseGPT (Frantar and Alistarh, 2023) formulates pruning as a layer-wise sparse regression problem, approximately solving it via a sequence of efficient Hessian updates and weight reconstructions. Wanda (Sun et al., 2024) computes the importance score of each weight based on the product of the magnitude of each weight and the norm of the corresponding input activation, identifying and removing weights with lower importance scores. OWL (Yin et al., 2024) identifies the correlation between pruning efficacy and the retention ratio of outliers, assigning different sparsity ratios to each layer based on the observed outlier ratio. RIA (Zhang et al., 2024b) introduces a metric that considers both weight and activation information, utilizing a permutation strategy for the input channels of weight matrices to enhance pruning performance. BESA (Xu et al., 2024) adopts a layer-wise pruning strategy, independently pruning each Transformer layer to minimize the reconstruction error between the outputs of pruned and dense Transformer layers, which avoids accumulating errors across layers.
Structured pruning.
Structured pruning focuses on broader network structures, such as neurons, attention heads, or even entire modules. LLM-Pruner (Ma et al., 2023) utilizes gradient information to identify interdependent structures within LLMs, pruning the least important groups and subsequently using Low-Rank Adaptation (LoRA) (Hu et al., 2022) to restore the performance of pruned models. LoRAPrune (Zhang et al., 2023) estimates the importance of pre-trained weights using LoRA gradients, iteratively removing redundant channels in the weight matrices and recovering the pruned models’ performance through fine-tuning. Sheared-LLaMA (Xia et al., 2024) learns a set of pruning masks to extract a sub-network with the specified target structure from the source model, employing a dynamic batch loading algorithm to adjust the data proportion of each domain based on the loss reduction rate in different domains. SliceGPT (Ashkboos et al., 2024) introduces the concept of computational invariance, achieving compression by removing rows or columns corresponding to smaller principal components in the weight matrix. LaCo (Yang et al., 2024) proposes a concise layer pruning approach, reducing model size by merging layers while maintaining the overall model structure. ShortGPT (Men et al., 2024) introduces a metric for measuring layer importance, achieving model compression by removing redundant layers.
Although unstructured pruning can maintain performance at higher pruning ratios, it often requires additional hardware or library support, making model acceleration impractical. Current structured pruning methods typically require retraining the model after pruning to avoid performance collapse. While layer pruning techniques like LaCo eliminate the need for additional retraining, their disregard for fine-grained block redundancy makes it challenging to avoid significant performance loss.
Concurrent and independent of our research, FINERCUT (Zhang et al., 2024a) also presents a fine-grained block pruning algorithm. However, their study does not delve into the rationale behind treating Transformer layers as two distinct sublayers for pruning purposes. In contrast, we began by conducting preliminary experiments that unveiled the fine-grained block redundancy within Transformer models. This discovery led us to propose the concept of minimal residual blocks. Furthermore, we explored how pruning various types of blocks affects the model’s performance. Furthermore, FINERCUT evaluates block importance by comparing the similarity between the output logits of the initial and pruned models. However, this metric may not adequately ensure that the pruned model generates coherent and semantically correct text, as it overlooks semantic information. In contrast, our approach measures block importance using the perplexity of the pruned model, which better reflects the fluency and quality of its output.
Figure 2: Illustration depicting that a Transformer layer can be subdivided into two residual blocks.
3 Methodology
The proposed fine-grained block pruning method (BlockPruner) is depicted in Figure 3. It begins by decomposing each Transformer layer into two minimal residual blocks (§3.1). We then evaluate the importance of each block using our proposed block importance metric (§3.2). Finally, we iteratively prune the block with the lowest importance (§3.3).
Figure 3: Overview of our BlockPruner. We iteratively calculate the importance score for each block (MHA or MLP) to obtain the block importance distribution, and subsequently remove the block with the lowest importance.
3.1 Minimal Residual Block
Most contemporary LLMs (Brown et al., 2020; Touvron et al., 2023a, b) are built upon the GPT architecture (Radford et al., 2019), which constitutes a decoder-only model comprising multiple Transformer layers, an embedding layer, and a language model head. As depicted in Figure 2, each Transformer layer can be decomposed into two primary residual blocks: the multi-head attention (MHA) block and the multi-layer perceptron (MLP) block.
Formally, consider the input hidden states of the $i$th Transformer layer, denoted as ${X}_{i-1}\in\mathbb{R}^{n\times d}$, where $n$ represents the length of the input sequence, and $d$ represents the hidden layer dimension of the model. The computational process within the $i$th Transformer layer can be represented as follows:
$X_{i}^{\prime}=\mathrm{MHA}(\mathrm{LN}(X_{i-1}))+X_{i-1},$ (1)
$X_{i}=\mathrm{MLP}(\mathrm{LN}(X_{i}^{\prime}))+X_{i}^{\prime}.$ (2)
Here, $\mathrm{LN}(\cdot)$ denotes the layer normalization module and $X_{i}^{\prime}\in\mathbb{R}^{n\times d}$ represents the intermediate hidden states after the MHA block.
Equations (1) and (2) indicate that both types of residual blocks can be abstracted into a same computational formula. Therefore, intuitively treating them as independent layers for pruning appears reasonable, which will be further validated by our subsequent experimental results.
3.2 Block Importance
While previous layer pruning methods (Men et al., 2024; Chen et al., 2024) rely solely on the similarity between layer inputs and outputs to measure layer importance, we contend that this approach considers only the local influence of the layer while neglecting its role in the overall model’s output. To address the drawback, we introduce perplexity as a measure of block importance. Specifically, we determine the importance score of each block by masking it and then computing the perplexity of the new model on a given dataset. Intuitively, a block with the lowest importance score indicates that its removal results in minimal performance degradation. This method more effectively captures each block’s overall impact on the model’s performance, thereby more accurately reflecting its significance.
Mathematically, perplexity is defined as the exponential of the average negative log-likelihood of a sequence of words. Given a sequence of words $w_{1},\ldots,w_{n}$ and a language model that predicts the probability $p_{\theta}(w_{i}|w_{<i})$ for each word $w_{i}$, the perplexity $\mathrm{PPL}$ is calculated as:
$\mathrm{PPL}=\mathrm{exp}(-\frac{1}{n}\sum_{i=1}^{n}\mathrm{\log}p_{\theta}(w_{i}|w_{<i})),$ (3)
where $p_{\theta}(w_{i}|w_{<i})$ denotes the probability of word $w_{i}$ given the preceding words in the sequence.
3.3 Iterative Search for Block Pruning
Unlike existing layer pruning techniques, which indiscriminately remove entire Transformer layers, we propose a novel fine-grained pruning strategy. This strategy selectively prunes multi-head attention (MHA) or multi-layer perceptron (MLP) blocks based on their defined importance. By employing this finer-grained pruning approach, we aim to better preserve the critical components and capabilities of the model while aggressively removing the less significant blocks.
For an LLM $\mathcal{M}$ with $L$ layers, we first divide them into $2L$ blocks, consisting of multi-head attention (MHA) and multi-layer perceptron (MLP) blocks. Then, we perform iterative pruning search on a calibration dataset $\mathcal{C}$ to sequentially prune $K$ blocks. The steps are outlined as follows:
Step 1: Mask Block. For each block $B_{i}$ (MHA or MLP) in $\mathcal{M}$, we generate a modified model $\hat{\mathcal{M}}$ by masking out this block.
Step 2: Calculate Importance. We compute the perplexity $P_{i}$ for the modified model $\hat{\mathcal{M}}$ on the calibration dataset $\mathcal{C}$ as the importance score for the masked block $B_{i}$.
Step 3: Sort and Prune. After computing the importance scores for all blocks, we sort these scores and remove the block with the lowest importance score from $\mathcal{M}$ to create a new model.
Step 4: Iterate. The aforementioned steps are iteratively repeated until $K$ blocks are removed.
Algorithm 1 Iterative Block Pruning
0:  Model $\mathcal{M}$ with $L$ layers, calibration dataset $\mathcal{C}$, number of blocks to remove $K$
0:  Pruned model $\mathcal{M}^{*}$
1:  $\mathcal{M}_{0}$leftarrow $\mathcal{M}$
2:  Split the model $\mathcal{M}_{0}$ into $2L$ blocks
3:  for $j=1$ to $K$ do
4:     for $i=1$ to $12L-j+$ 1 do
5:        Create model $\hat{\mathcal{M}}$ by masking block $B_{i}$;
6:        Compute the perplexity $P_{i}$ for $\hat{\mathcal{M}}$ on the calibration dataset $\mathcal{C}$;
7:     end for
8:     Sort the blocks based on their perplexities;
9:     Remove the block with the lowest perplexity from $\mathcal{M}_{j-1}$ and obtain $\mathcal{M}_{j}$;
10:  end for
11:  $\mathcal{M}^{*}$leftarrow $\mathcal{M}_{K}$
12:  return  Pruned model $\mathcal{M}^{*}$
By iteratively removing the blocks with the lowest importance scores, we aim to prune the LLM while minimizing performance degradation on the calibration dataset $\mathcal{C}$. This fine-grained block pruning approach provides a more targeted method for pruning LLMs compared to traditional layer-level pruning techniques, thereby facilitating more efficient model compression while better preserving the model’s performance. The detailed procedure for this pruning process is outlined in Algorithm 1.
4 Experiments
In this section, we first introduce the experimental setups and then present the main results.
4.1 Experimental Setups
Models.
To validate the widespread effectiveness of our pruning method, we experiment with three series of models: Llama2 (Touvron et al., 2023b), Baichuan2 (Yang et al., 2023), and Qwen1.5 (Bai et al., 2023). These models share analogous architectures as described in equations (1) and (2). Due to computational constraints, we employ 7B and 13B models for Llama2 and Baichuan2, respectively, and 7B and 14B models for Qwen1.5.
Baselines.
Model Method Ratio (%) PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg. Score
Llama2-7B Dense 0 5.47 79.05 69.06 75.99 74.54 46.16 68.96
SliceGPT 21.45 30.74 72.42 59.91 56.04 63.64 37.12 57.83
LaCo 21.02 50.39 68.34 60.46 54.08 55.39 35.84 54.82
RM 21.02 676.80 54.46 49.25 29.22 34.43 22.53 37.98
ShortGPT 21.02 18.45 70.24 65.90 62.63 56.06 36.09 58.18
BlockPruner 21.99 11.51 74.21 62.43 65.87 61.07 37.29 60.17
Llama2-13B Dense 0 4.89 80.52 72.14 79.36 77.36 49.23 71.72
SliceGPT 21.52 23.95 74.32 65.59 60.71 68.52 42.41 62.31
LaCo 24.37 13.97 72.42 59.27 60.44 54.34 34.56 56.21
RM 24.37 10.08 73.72 66.61 66.80 66.12 41.98 63.05
ShortGPT 24.37 20.06 72.74 70.80 67.80 60.35 41.30 62.60
BlockPruner 25.12 8.16 76.93 66.30 72.20 65.82 41.38 64.53
Baichuan2-7B Dense 0 6.04 77.48 68.27 72.18 72.98 42.75 66.73
LaCo 21.57 26.46 68.28 58.56 51.50 52.90 28.50 51.95
RM 21.57 189.78 59.96 52.33 30.87 38.17 23.63 40.99
ShortGPT 21.57 31.05 63.71 62.67 50.01 47.31 30.72 50.88
BlockPruner 22.45 15.38 69.75 61.48 58.09 58.08 33.02 56.08
Baichuan2-13B Dense 0 6.66 78.84 70.40 75.23 74.07 47.70 69.25
LaCo 22.68 27.07 70.89 58.01 54.00 57.11 32.94 54.59
RM 22.68 17.70 68.99 67.88 63.78 57.49 37.54 59.14
ShortGPT 22.68 20.69 69.31 68.27 61.71 56.52 36.69 58.50
BlockPruner 24.19 15.36 71.44 64.01 64.20 59.81 37.88 59.47
Qwen1.5-7B Dense 0 7.95 79.22 66.46 76.92 62.16 42.66 65.48
LaCo 20.97 39.23 70.40 58.64 56.35 46.89 32.85 53.03
RM 20.97 2026.31 67.36 49.88 42.00 54.17 28.58 48.40
ShortGPT 20.97 49.88 69.53 62.12 58.87 43.60 32.17 53.26
BlockPruner 21.83 20.58 71.71 55.56 59.31 53.70 33.28 54.71
Qwen1.5-14B Dense 0 7.44 79.87 70.56 79.41 68.48 47.01 69.07
LaCo 22.25 16.32 71.55 58.33 60.16 53.70 34.04 55.56
RM 22.25 55.99 67.08 53.28 42.08 50.72 29.01 48.43
ShortGPT 22.25 1237.21 58.60 55.96 36.16 38.09 34.81 44.72
BlockPruner 23.72 15.67 75.24 61.48 66.92 59.51 39.08 60.45
Table 1: Zero-shot downstream task performance of various models using different pruning methods. “Dense” represents the original, unpruned models. “PPL” means the perplexity on Wikitext2. All evaluations are conducted using the same configuration to ensure comparability.
We compare our method with several state-of-the-art structured pruning methods. The specific baseline methods include SliceGPT (Ashkboos et al., 2024), LaCo (Yang et al., 2024), ShortGPT (Men et al., 2024), and Relative Magnitude (Samragh et al., 2023; Men et al., 2024). SliceGPT achieves pruning by removing rows or columns corresponding to smaller principal components in the weight matrix. LaCo merges model layers from deep to shallow, using model output representations to calculate thresholds to avoid over-merging. ShortGPT eliminates redundant layers by calculating Block Influence. Relative Magnitude (RM) uses $||\frac{f(x)}{x+f(x)}||$ as an importance metric for layers, where $f(.)$ represents the non-residual part of the Transformer layer, and employs the same pruning method as ShortGPT. For SliceGPT, we used the official implementation [^2] As SliceGPT’s official code does not support Baichuan2 and Qwen1.5, we only employ it on the Llama2 series models.. For LaCo, we implemented it based on their code and controlled the number of pruned layers by adjusting the merging threshold. For ShortGPT and RM, we reproduced the results based on their paper descriptions.
Data and GPUs.
In our main experiment, we utilize the Alpaca dataset (Taori et al., 2023) to calculate importance scores. For our method, we employ only 256 samples to compute perplexity, and we discuss the influence of varying sample sizes in Section 5.4. Moreover, we observe that ShortGPT and Relative Magnitude are not sensitive to different numbers of samples. The effect of sample size on ShortGPT and Relative Magnitude is detailed in Appendix H. Nevertheless, we used the same number of samples as our method for consistency. All experiments in this study are conducted using two RTX 4090 GPUs.
Evaluations.
Following SliceGPT, we use LM Evaluation Harness (Gao et al., 2023) for evaluation and validation on five well-known benchmarks: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018). We also utilize Wikitext2 dataset (Merity et al., 2016) for evaluating the perplexity after pruning. More comprehensive details of can be found in Appendix C.
4.2 Main Results
Previous structured pruning methods typically prune less than 30% of the parameters. Therefore, in our main experiments, we controlled the pruning ratio within this range. Since it is challenging to achieve identical pruning ratios across different methods and models, we select the closest available pruning ratios for comparison. The experimental results are presented in Table 1.
As shown in the results, our BlockPruner method significantly outperforms previous structured pruning baselines in terms of average performance and achieves the best results across most benchmarks, even though the pruning ratios in our method are slightly higher than that of baselines. We also observe that Llama2-13B maintains better performance at higher pruning ratios compared to Llama2-7B, with Baichuan2 and Qwen1.5 exhibiting similar behavior. This suggests that as the model scale grows, so does the number of redundant blocks, allowing for more pruning space.
Furthermore, it’s noteworthy that models exhibiting lower perplexity on the Wikitext2 dataset generally outperform those with higher perplexity on the same dataset. This underscores the potential of perplexity as a metric reflecting model performance. Notably, despite our method conducting pruning searches on the Alpaca dataset, it achieves lower perplexity on the Wikitext2 dataset.
Finally, we observe that while approaches such as ShortGPT and Relative Magnitude result in a significant decline in model performance across different tasks, BlockPruner stands out by avoiding such drastic reductions. This suggests that our proposed block pruning method effectively mitigates performance degradation during the pruning process. Due to space constraints, we have moved the details of pruning baselines and comparisons across various pruning ratios to Appendix I.
5 Analyses
5.1 Ablation Study
To assess the influence of various key operations within the proposed pruning algorithm on its performance, we undertake a thorough ablation study across six models. In particular, we first drop the iterative search procedure and directly remove blocks with the lowest importance scores. Then, we substitute the fine-grained block pruning with a coarser-grained layer pruning approach. The results of these experiments are shown in Table 2.
The experimental findings highlight that solely relying on the perplexity metric without incorporating a search component can result in subpar pruning results and even performance deterioration. This phenomenon may stem from the intrinsic nature of perplexity, which, unlike other importance metrics focusing solely on local block influence, is inherently influenced by the interaction among multiple blocks due to its derivation from the model’s output calculation. While perplexity aids in identifying redundant blocks within the model, it doesn’t directly yield an optimal pruning sequence.
Furthermore, pruning at the layer level rather than the fine-grained block level yields less robust performance. This observation indicates that the model contains fine-grained redundancies, and segmenting layers into smaller blocks for pruning allows for more efficient removal of this redundancy, thereby better preserving the model’s capabilities.
Model Method Ratio (%) Avg. Score
Llama2-7B BlockPruner 21.99 60.17
- search 20.95 55.89 (-7.11%)
- block 21.02 58.63 (-2.56%)
Llama-2-13B BlockPruner 25.12 64.53
- search 25.08 58.58 (-9.21%)
- block 24.37 62.91 (-2.51%)
Baichuan2-7B BlockPruner 22.45 56.08
- search 22.39 38.81 (-30.80%)
- block 21.57 54.76 (-2.36%)
Baichuan2-13B BlockPruner 24.19 59.47
- search 24.19 55.95 (-5.92%)
- block 24.95 58.22 (-2.10%)
Qwen1.5-7B BlockPruner 21.83 54.71
- search 20.90 37.72 (-31.06%)
- block 20.97 52.66 (-3.75%)
Qwen1.5-14B BlockPruner 23.72 60.45
- search 22.98 40.80 (-32.51%)
- block 22.25 60.10 (-0.58%)
Table 2: Average score of ablation study of BlockPruner on downstream tasks. “- search” indicates dropping the iterative search procedure and directly removing blocks with the lowest importance score. “- block” means we substitute the fine-grained block pruning with a coarser-grained layer pruning approach.
5.2 Redundancies Between MHA and MLP
To investigate the significance and roles of the MHA and MLP modules in modern LLMs, we conduct pruning experiments focusing exclusively on MHA or MLP blocks. We apply this pruning strategy to two models of varying sizes, Llama2-7B and Llama2-13B, while keeping the pruning ratios below 33%. The results illustrated in Figure 4 reveal several notable observations.
Before reaching a pruning ratio of 17%, pruning only the MHA blocks results in less performance loss compared to pruning MLP blocks and even matches the performance of mixed pruning. This indicates that MHA modules in LLMs may possess greater redundancy than initially anticipated, whereas MLP modules are relatively less redundant. However, when the pruning ratio surpasses 17%, further pruning of MHA blocks leads to a sharp decline in performance. This trend suggests that as pruning advances, the redundant MHA blocks are progressively removed, leaving only the crucial MHA blocks. Moreover, in the larger model, the sharp decline in performance occurs at higher pruning ratios. This observation is consistent with our previous findings, suggesting that larger models contain more redundant blocks. Such redundancy may stem from factors like insufficient training, resulting in higher initial redundancy.
Figure 4: The impact of pruning MHA and MLP individually on model performance. “MHA&MLP” represents the original BlockPruner algorithm.
We also examine the proportion of MHA blocks removed during pruning. Specifically, we present the number of MHA and MLP blocks removed at different pruning stages. In Figure 5 (left), we set the number of removed blocks to 60. In Figure 5 (right), the models have 22 and 28 blocks removed, respectively, maintaining a pruning ratio of 30%.
The results in Figure 5 (left) for both models reveal a consistent tendency to initially remove only MHA blocks. As the pruning process progresses and more blocks are removed, the proportion of MHA blocks being pruned follows a zigzag downward trend. Notably, the curve for Llama2-13B shifts to the right compared to Llama2-7B, suggesting that the larger model contains more redundant MHA blocks. This is further emphasized in Figure 5 (right), where, at the same pruning ratio, Llama2-13B prunes more MHA blocks than Llama2-7B. Additionally, given that our pruning method tends to remove more MHA blocks at equivalent pruning ratios, it can significantly reduce the usage of the key-value (KV) cache (Pope et al., 2023) in MHA, which potentially accelerate the inference process. To validate this, we also conducted a comparison of the inference speed among various models obtained through different pruning methods, with the results detailed in Appendix E.
Figure 5: Left: The proportion of MHA blocks removed during the pruning process, relative to the total number of removed blocks. Right: The number of different blocks removed from models at a pruning ratio of 30%.
5.3 Perplexity for Block Redundancy
In this section, we explore the impact of different block importance metrics. Generally, Block Influence (BI) and Relative Magnitude (RM) measure the importance of a block based solely on its input and output hidden states, thereby reflecting the block’s local influence. In contrast, perplexity is derived from the model’s output representations and thus can better measure a block’s overall influence.
However, as indicated in the ablation study, using perplexity without the iterative search procedure leads to a significant decline in performance. This suggests that perplexity alone is not an effective block importance metric. Instead, it is better suited for dynamic pruning algorithms that offer greater flexibility compared to static algorithms.
As illustrated in Figure 6, when BI and RM are applied in dynamic pruning algorithms, they sometimes achieve performance comparable to perplexity at lower pruning ratios. However, as the pruning ratio increases, their limitations become evident, resulting in a sharp decline in model performance. This suggests that these local metrics do not adequately capture the impact of different blocks on the model’s overall performance.
In summary, perplexity leverages global information to effectively measure block redundancy, especially when used with a dynamic pruning strategy. This combination captures the complex interactions among blocks. In contrast, local metrics like Block Importance (BI) and Relative Magnitude (RM) are useful in specific scenarios but don’t reflect the overall contribution of blocks to the model, particularly at higher pruning ratios.
Figure 6: The impact of different block importance metrics on the pruning performance of BlockPruner
5.4 Impact of Data on Pruning
In the work on SliceGPT (Ashkboos et al., 2024), the authors also used the Wikitext2 (Merity et al., 2016) and Alpaca (Taori et al., 2023) datasets for pruning experiments. They observed that the Alpaca dataset often yielded better pruning results. In our study, we obtain similar findings. As shown in Figure 7 (left), when pruning Llama2-7B, the performance across different pruning ratios is significantly higher when using the Alpaca dataset compared to Wikitext2. We hypothesize that this may be due to the Alpaca dataset being an instruction-following dataset, which is more closely aligned with downstream tasks. This suggests that the choice of dataset has a significant impact on the final pruning performance of the model.
To determine the appropriate sample size and analyze its impact on the pruning performance of BlockPruner, we extract varying numbers of instances from the Alpaca dataset and conduct pruning experiments using Llama2-7B. The results presented in Figure 7 (right) indicate that increasing the sample size beyond 256 yields no significant improvement in the pruning effect of BlockPruner. Therefore, we set the number of samples to 256.
Figure 7: Left: The performance of BlockPruner on the Alpaca and Wikitext2 datasets using a calibration dataset of 256 samples. Right: Impact of sample sizes on BlockPruner’s performance on Alpaca, with the numbers indicating the sample sizes used.
6 Conclusion
In this work, we introduce BlockPruner, a novel structured pruning approach for efficiently pruning large language models (LLMs). BlockPruner decomposes Transformer layers into two minimal residual blocks and employs a new block importance metric along with a pruning search algorithm to iteratively remove redundant blocks. Extensive experiments across various models show that our method outperforms other baselines in post-pruning performance. Our findings highlight the presence of fine-grained block redundancy in LLMs and reveal significant differences in redundancy levels among different block types. We hope our work contributes to a deeper understanding of the importance of various blocks within LLMs.
7 Limitations
Our current work has three potential limitations. First, while perplexity serves as a useful indicator of block importance, it may not be the optimal metric. Second, while our proposed pruning search algorithm is effective, other combinatorial optimization algorithms might identify superior pruning sequences. Lastly, due to constraints in computational resources, we did not apply our method to prune larger models. Nevertheless, our approach is highly scalable and readily adaptable for pruning larger models in future research.
Appendix A Details of Implementations
In this section, we detail our experimental setup. We sampled from the Alpaca dataset with a fixed random seed of 42. For SliceGPT, we followed the original paper’s configuration, using 1024 samples, a sparsity ratio set at 30%, and a maximum sequence length of 2048. For ShortGPT, RM, and BlockPruner, we sampled 256 samples from the dataset, with the same maximum sequence length of 2048. For LaCo, we adjusted the merging threshold using the provided code and data to achieve the corresponding pruning ratio.
Appendix B Details of Datasets
B.1 Pruning Datasets
Alpaca
is a general instruction-following dataset containing 52,000 questions. Each sample comprises three fields: instruction, input, and response. We selected 10% of the dataset and utilized 256 samples for the main experiments. Perplexity calculation was performed uniformly across all text in the samples without differentiation between fields.
B.2 Evaluation Datasets
All downstream task datasets were partitioned and evaluated using the default configuration of LM Evaluation Harness.
Wikitext-2
is a collection of over 100 million tokens extracted from verified Good and Featured articles on Wikipedia. This dataset is commonly used to measure the quality of a model’s text generation. We employed samples from the pre-split test set for calculating perplexity.
PIQA
is a dataset designed to evaluate natural language models’ understanding of physical commonsense. It employs a multiple-choice format where, given a goal and two possible solutions, the model must select the most appropriate solution.
WinoGrande
is an extensive dataset to evaluate models’ commonsense reasoning capabilities. It comprises 44,000 questions. The dataset features fill-in-the-blank tasks with binary options, aiming to select the correct option for a given sentence that requires commonsense reasoning.
HellaSwag
is also a dataset designed to assess models’ commonsense reasoning abilities, specifically to highlight the limitations of current models in handling commonsense natural language reasoning tasks. Despite being trivial for humans (with >95% accuracy), the dataset presents significant difficulties for models. The evaluation is conducted using four-way multiple-choice questions.
ARC
dataset comprises 7,787 multiple-choice science exam questions sourced from various origins. Each question typically offers four answer options. These questions are categorized into two distinct difficulty sets: 2,590 questions for Challenge Set and 5,197 for Easy Set.
Appendix C Details of Evaluations
To ensure a fair and comprehensive comparison, we employed the same version of the LM Evaluation Harness as used in the SliceGPT experiments and obtained evaluation scores under identical experimental configurations. These scores closely match those reported in the SliceGPT paper, as detailed in Table 3. For consistency, we present our reproduced results in the main experiments.
Model Method Ratio(%) Avg.Score
Llama2-7B $\mathrm{SliceGPT}$ 21.45 57.93
$\mathrm{SliceGPT}^{*}$ 21.45 57.83
Llama2-13B $\mathrm{SliceGPT}$ 21.52 62.34
$\mathrm{SliceGPT}^{*}$ 21.52 62.31
Table 3: Comparison of average performance on downstream tasks between the official SliceGPT results and our reproduced results (indicated by “*” for our results).
To evaluate the performance of pruned models on downstream tasks, we utilized five multiple-choice QA datasets: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c. Additionally, to assess text generation quality, we calculated perplexity using the test set of the Wikitext2 dataset. For the downstream task evaluations, we adhered to the default evaluation parameters and zero-shot settings, with a batch size set to 1. For perplexity calculations, the maximum text length was set to 2048, maintaining a batch size of 1 as well.
Appendix D Ablation Experiments under Higher Sparsity
The motivation behind BlockPruner is to maintain model performance more effectively through fine-grained block pruning. Evaluating the performance differences of block pruning at varying granularities under higher sparsity is essential, as demonstrating effectiveness under such conditions would further corroborate our motivations and claims. In light of this, we conducted ablation experiments with higher sparsity ratios on LLaMA-2-7B and LLaMA-2-13B models and found that our motivation and claims remain valid. The results are presented in Table 4.
Model Unit Ratio(%) PPL($\downarrow$) Avg.Score
Llama2-7B Block 30.00/37.02/43.03 16.28/27.47/49.65 55.02/49.36/46.95
Layer 30.03/36.04/42.05 16.58/27.05/60.85 53.04/47.32/43.91
Llama2-13B Block 31.62/36.52/41.39 9.64/12.54/17.02 61.52/59.34/54.15
Layer 31.68/36.56/41.43 10.48/13.29/18.04 59.31/56.17/51.61
Table 4: Average score of BlockPruner at different pruning granularities under higher sparsity.
Appendix E Inference Speed after Pruning
Since our method reduces more MHA blocks at the same pruning ratio compared to other methods, it consequently reduces the usage of the KV cache at equivalent pruning ratios. This reduction can potentially accelerate the inference speed of LLMs. To verify that our method indeed accelerates the model’s inference process, we compared the time taken to generate 128 tokens by models obtained from different pruning methods with KV caches. To ensure the robustness of the experimental results, we repeated each configuration 20 times and took the average inference time as the result. As observed from the Table 5, our method achieves the highest speedup at comparable pruning ratios.
Model Method Ratio(%) Inference Time (ms) Speedup
Llama2-7B Original 0.00 4044.30 1.00
BlockPruner 24.00 2747.88 1.47
SliceGPT 24.47 3226.68 1.25
ShortGPT 24.03 3094.36 1.31
Llama2-13B Original 0.00 7285.73 1.00
BlockPruner 21.05 3873.20 1.88
SliceGPT 21.52 4099.08 1.78
ShortGPT 21.93 4111.65 1.77
Table 5: The inference speed differences among models obtained using different pruning methods, where “Original” denotes the unpruned model.
Appendix F Time Costs of Various Pruning Methods
Our approach relies on PPL to determine block importance, which requires calculating PPL before pruning, making it challenging to design a more efficient pruning strategy. We have compared other block importance metrics (in Section 5.3) but found that PPL still preserves the model’s performance best. Moreover, since our method better maintains model performance and pruning is one-time without increasing subsequent inference overhead, so we believe the trade-off is worthwhile. The comparison results of pruning times between BlockPruner and other methods are presented in Table 6.
Model BlockPruner SliceGPT ShortGPT RM LaCo
Llama2-7B 45 mins 2 hours 9 mins 2 mins < 1 mins 2 mins
Llama2-13B 2 hours 27 mins 3 hours 30 mins 2 mins < 1 mins 24 mins
Table 6: Execution time of BlockPruner and other pruning methods in the main experiment.
Appendix G Post-training after Pruning
We sampled 8,000 instances from the Alpaca dataset and performed simple post-training on the pruned Llama2-7B and Llama2-13B obtained through BlockPruner using LoRA. We trained all linear layers except for the embedding layer and the language model head. The parameters LoRA rank and LoRA $\alpha$ were set to 32 and 10, respectively, with a learning rate of 2e-4 and a training batch size of 1. We also config the gradient accumulation steps to 4 and use a linear learning rate scheduler. We controlled the pruning ratios within the range of 24% to 33%. The results are shown in Figure 8. It can be seen that after training, our models showed further improvement at different pruning ratios. The Llama2-7B and Llama2-13B models recovered to 89% and 92% of the performance of the unpruned models, respectively, when pruned by approximately 1/4.
Figure 8: Average score of BlockPruner with varying pruning ratios before and after post-training.
Appendix H Sensitivity to Sample Size
ShortGPT uses Block Influence as the importance metric for layers, while RM uses Relative Magnitude. The former calculates the similarity between the input and output hidden states of a layer, while the latter utilizes the input and the non-residual part of the output. In our preliminary experiments, we found that these two metrics are not sensitive to sample size. We sampled different numbers of instances from the test set of the Alpaca dataset to observe their impact on these metrics, and the results are shown in Figure 9. We can see that all the lines almost overlap, indicating that Block Influence and Relative Magnitude are not sensitive to the sample size. We speculate that this may be due to the limited information provided by the changes in the input and output of a single layer.
Figure 9: The changes in Block Influence and Relative Magnitude of the model under different sample sizes.
Appendix I Varying Pruning Ratios
To broadly demonstrate the superiority of our method, we present the pruning effects of BlockPruner, ShortGPT, and Relative Magnitude on six representative large models at different pruning ratios. As depicted in Figure 10, our method effectively minimizes performance loss throughout the pruning process, avoiding any sudden drops in performance. In contrast, RM exhibits significant instability and is prone to performance collapse. ShortGPT performs relatively well, but in the pruning experiments on Qwen1.5-14B, it also leads to severe performance degradation at higher pruning ratios. Overall, the advantages of our method become more pronounced as both model size and pruning ratio increase.
Figure 10: Average score of BlockPruner with varying pruning ratios compared with ShortGPT and RM.